OpenStudy (anonymous):

An object is thrown directly up (positive direction) with a velocity (vo) of 20.0 m/s and do= 0. How high does it rise (v = 0 cm/s at top of rise). Remember, acceleration is -9.80 m/s2.

OpenStudy (anonymous):

AS the object is thrown with some velocity therefore, initial velocity(u) would be zero and final velocity is 20 m/s In order to determine how high does it rise we will use third eqn of motion , that is, \[v ^{2} - u ^{2} = 2*g*h\] where g is acceleration due to gravity

OpenStudy (anonymous):

you can also think of it in terms of conservation of energy. so initially, all energy of the ball(relative to earth) is its kinetic energy(due to initial velocity) finally at the top, all energy is in the form of extra potential energy(due to gained height) hence, (mu^2)/2 = mgh therefore h = u^2/2g = (20*20)/(2*9.8) = 20.408 m

OpenStudy (anonymous):

thanks